What is the perfect standard normal distribution?
The standard normal distribution is a subset of a normal distribution. It has the properties of mean equal to zero and a standard deviation equal to one. There is only one standard normal distribution and no others so it could be considered the "perfect" one.
About 63.3968%
Right. Winning AT LEAST once out of 100 tries is equal to NOT(NEVER winning in one of 100 tries). The one in the brackets is 0.99 ^100, which is about 0.33. Taking the NOT is subtracting that result from one, yielding about 0.63396766. To get a percentage multiply this by 100.
The answer is 8.
8+32=40 and 8x5=40 ****** thx. =]
What is the theoretical probability of tossing three coins?
it is 1/3
uhhh be specific because on apex its 3/8
The probability of any event is always a number from 0 to 1, inclusive. It can never be 12.
What is the probability of a meteor falling on your house?
hmmm...check this out: #houses in city X #citys in province/state X # provinces in country X #countrys in hte world. the answer is the probability X equals times (as in math.).
What should occur in a probability experiment?
An event, unless it already had been occured and the experiment tries to resolve posterior probabilities on the event
What is the definition of mutually inclusive events?
Any two events in which one cannot happen without the other.
How much does a hydrogen bomb cost?
Enough in which the president would be the one to ask you at an immensely dire time. In English terms, only if we cannot end a war unless we use a thermonuke.
If A and B are independent events then are A and B' independent?
if P(A)>0 then P(B'|A)=1-P(B|A)
so P(A intersect B')=P(A)P(B'|A)=P(A)[1-P(B|A)]
=P(A)[1-P(B)]
=P(A)P(B')
the definition of independent events is if P(A intersect B')=P(A)P(B')
that is the proof
The probability is 9/1 that the two people are not Norwegians.
Probability of drawing a blue marble first is 4 in 8 (or 50%)
Probability of drawing a blue marble second is 3 in 7 (or 42.85714%)
Probablility of drawing blue then blue is the two above multiplied 0.5 * 0.4285714
Which is 0.212142407 or 21% or One in Five.
How is probability used in sports?
Probability could be used in a sport to test or experiment the chances of getting an injury or losing/winning a game.
There are 26 red cards in a standard deck, 13 hearts and 13 diamonds.
That's excluding any jokers.
How do you say Books in french?
"Un livre" (lee-vru) is the word for a paper book.
To book, for example, a hotel room you would use "reserver" -
I wish to book a room is "Je veux réserver une chambre."